OpenStudy (anonymous):

anti derivative of 3/x?

OpenStudy (anonymous):

=integral 3*1/x = 3 ln x

OpenStudy (anonymous):

actually 3 ln |x|

OpenStudy (anonymous):

so with or without absolute sign?

OpenStudy (anonymous):

with

OpenStudy (anonymous):

can any of you help me onsome differential equation problems?

OpenStudy (anonymous):

idk but ill give it a shot

OpenStudy (anonymous):

solve the initial value problem of y'=xe^y-2e^y if y(0)=0

OpenStudy (anonymous):

already confused sorry cant help you

OpenStudy (anonymous):

thats ok

OpenStudy (anonymous):

\[y'=xe^y-2e^y \] \[y'=e^y(x-2)\] \[\frac{dy}{dx}=e^y(x-2)\] \[dy=e^y(x-2)dx\] \[\frac{dy}{e^y}=(x-2)dx\] Integrate both sides! :)

OpenStudy (anonymous):

\[\int\limits_{}^{}e^{-y} dy =\int\limits_{}^{}(x-2)dx\]

OpenStudy (anonymous):

\[-e^{-y}=\frac{x^2}{2}-2x+C\]

OpenStudy (anonymous):

Now apply initial condition to find C

OpenStudy (anonymous):

replace x with 0 and replace y with 0 and solve for C